# What’s in a Name? Auditing Large Language Models for Race and Gender Bias

### v2

@@ Abstract
We employ an audit design to investigate biases in state-of-the-art large language models, including GPT-4. In our study, we prompt the models for advice involving a named individual across a variety of scenarios, such as during car purchase negotiations or election outcome predictions. We find that the advice systematically disadvantages names that are commonly associated with racial minorities and women. Names associated with Black women receive the least advantageous outcomes. The biases are consistent across 42 prompt templates and several models, indicating a systemic issue rather than isolated incidents. While providing numerical, decision-relevant anchors in the prompt can successfully counteract the biases, qualitative details have inconsistent effects and may even increase disparities. Our findings underscore the importance of conducting audits at the point of LLM deployment and implementation to mitigate their potential for harm against marginalized communities.
@@ 1 Introduction
Large Language Models (LLM) have dramatically surged in popularity over the recent years. Since the release of ChatGPT, LLMs - especially those with an accessible chat interface - have not only been used by experts, but are also becoming an increasingly common tool with significant benefits for laypeople. To that end, many commercial actors have already begun implementing LLMs in their operations, ranging from customer-facing chatbots to internal decision support systems [14, 6].
@@
The fairness of AI algorithms, including LLMs, has been a pernicious issue, motivating a growing literature and community of AI ethics research [8]. Disparities across gender and race, among other attributes, have especially preoccupied this field [4], leading to efforts to include bias auditing as an important component of AI harm mitigation in policy discussions and regulatory frameworks [28].
@@
Mitigating biases arising from the explicit use of race or gender in the prompt is comparatively straightforward. For instance, popular models like GPT-4 often refuse to provide an answer when prompted to produce information about a hypothetical individual when given that individual’s race. Similarly, companies that have access to sensitive features of their customers may simply foreclose access to this information from the LLM.
@@
However, biases can materialize not only through the explicit use of sensitive characteristics, but also by utilizing features that are (strongly) correlated with a person’s protected attributes. The impact of these features can be harder to mitigate, because (1) their potential to cause disparate outcomes is often less salient, and (2) these features may contain information that otherwise improves the utility of the model. Among the potential candidates with these characteristics, we focus on an individual’s name as a feature of particular pertinence. Names strongly correlate with perceptions of race, raising the risk of creating significant disparities in model outputs, in turn harming marginalized communities. At the same time, in many practical applications, removing names might only come at a substantial cost. For instance, a chatbot that directly interacts with customers might significantly improve the experience via personalization if given access to the user’s name.
@@
In this study, we assess the name-sensitivity of the output produced by state-of-the-art language models. The names we choose are perceived to strongly correlate with race and/or gender, and we use direct model prompting as input to the models. Our assessment encompasses 42 idiosyncratic prompts. These prompts approximate use cases for 14 domains in which language models could be deployed to give advice to laypeople, such as in negotiations over the purchase of a car or in predicting election outcomes. We also assess how name-sensitivity interacts with the level of other useful information the model has access to in generating its output.
@@
We find significant disparities across names associated with race and gender in most scenarios we investigate, with varying effect sizes. The results are qualitatively similar across different models, including GPT-4, GPT-3.5 and PaLM-2. Overall, we find that names associated with white men yield the most beneficial predictions, while those associated with Black women generate outcomes that disadvantage the individual in question. Providing the model with qualitative context about the person has an inconsistent effect on biases, at times amplifying and at times decreasing observed disparities; while a numeric anchor effectively removes name-based disparities in most scenarios we investigate. Our findings also suggest that the observed disparities are the result of a systematic bias, rather than the result of a few name outliers.
@@
Overall, the results suggest that the model implicitly encodes common stereotypes, which in turn affects the model response. Because these stereotypes typically disadvantage the marginalized group, the advice given by the model does as well. Our findings suggest name-based differences commonly materialize into disparities to the disadvantage of women, Black communities, and in particular Black women. The biases are consistent with common stereotypes prevalent in the U.S. population.
@@
The findings show that despite efforts to mitigate biases and mount guardrails against disparate association with sensitive characteristics such as race and gender, LLMs encode biases that translate into disparate outcomes. Despite earlier concerns over bias, even the latest models, such as GPT-4, are not immune to this problem. Moreover, we show these name-based biases translate into disparate quantitative advice, effectively suggesting less favorable treatment of women and Black people. The findings raise concerns for businesses that seek to incorporate LLMs into their operations, suggesting that masking race and gender may not be enough to prevent unwanted disparities. The findings show that bias is pervasive and highlight the need for audits at the point of deployment and implementation, and not only at the development phase.
@@ 2.1 Defining Bias in Audit Studies
U.S. anti-discrimination laws generally encompass two distinct types of discriminatory conduct. First, there is “disparate treatment”, which refers to policies or actions that are conditioned on protected characteristics like race or gender. As such, disparate treatment corresponds to common intuitive understandings of discrimination in which an individual receives a certain cost or benefit because of their race/gender. Disparate treatment by governmental actors is scrutinized and generally outlawed under the Fourteenth Amendment of the U.S. Constitution, and is similarly illegal in most decision-making by private actors due to the existence of several federal and state laws.
@@
Connecting this legal framework to audit studies, it becomes apparent that audit designs are not directed at assessing bias in the form of disparate treatment. This is because, while they identify the causal effect of a feature strongly correlated with race, most audit studies do not directly identify the impact of race. [^1] Some legal scholars disagree with that conclusion on normative grounds, see [24]. Instead, our audit study identifies the impact of names on the output of a language model. But because names strongly correlate with race/gender, any disparities we observe may constitute bias in the form of disparate impact. To make that determination conclusively, it would be required to examine whether the disparities are justified, an assessment that will vary with the individual context.
@@ 2.2 Prior Literature
There is a substantial literature assessing bias in algorithms, including in medicine and health care [21, 23, 26, 11], law [13, 2, 5, 16, 12, 10, 31, 20], and education [1, 17]. The associated field is also referred to as “algorithmic fairness” [8], and its primary focus lies on assessing potential biases in algorithms that are used to assist human decision making. Researchers have also examined biases in automated speech recognition systems [18] and facial recognition systems [15], among others.
This study focuses on biases in language models. Previous attempts to detect such biases follow a variety of different methodologies. One common approach seeks to highlight implicit associations in the internal model representation of sensitive categories (like race or gender) and other desirable or undesirable traits or objects. An early example of this approach applied to word embedding models like word2vec [22] is the Word Embedding Association Test (WEAT) as introduced by Caliskan et al [4]. Under this test, the embedding representations of words representing sensitive attributes (like race or gender categories) are compared to the embeddings of a target vocabulary such as that formed by the Implicit Association Test (IAT). With the advent of more complex large language models, this approach has been adapted to exploit the relationship between references and objects in sentences where that relationship is ambiguous. For instance, Kotek et al [19] query a variety of LLMs with sentences such as ”the doctor phoned the nurse because she was late” asking the model to state who was late; and Sheng et al use completion for sentences such as ”the man worked as” to measure the regard a model has for a certain gender or racial/ethnic group [27]. However, implicit associations only represent one way in which biases can manifest. In addition, relying on implicit associations for the identification of biases may not represent an approach that is easily amenable to the different contexts in which the deployment of LLMs is contemplated. For instance, when language models provide negotiation advice, there may only be loose relationship between biases arising from implicit associations and those that substantively affect the negotiation strategy.
@@
Audit studies have a long tradition in assessing biases in human decisions, going back to the civil rights movement [28]. Historically, they have involved pairs of ”testers” who go through the process of seeking benefits such as employment or housing. The pairs were made to look and behave similarly, with the main difference that a sensitive attribute–like race or gender–differs across the individuals in the pair. By measuring differences in outcomes, the researchers could identify biases in the decision making process of the entity under investigation (e.g., a housing corporation) as they relate to the sensitive attribute [32, 25].
One particularly well-known example of an audit analysis is the resume correspondence study first conducted by Bertrand and Mullainathan [3]. The authors studied bias in hiring by submitting resumes to job postings, varying only the name of the applicant. The authors used stereotypical African-American, White, Male, and Female names as proxies for race and gender. The study has become a particularly popular example of auditing and has been replicated several times with variations, including in the audit of LLMs.
For example, Veldanda et al. [29] task LLMs (GPT-3.5, Bard, Claude and Llama) with matching resumes to job categories. They find no evidence of bias across race and gender, although the models displayed biases in regards to pregnancy status and political affiliation. In contrast, Wan et al.[30] task popular LLMs (GPT-3.5 and Alpaca) with crafting reference letters based on biographical details. They find substantial gender biases along the lexical content and language style the models output.
@@
Second, we use quantitative and continuous or granular discrete outcomes (see section 3.1), unlike previous efforts which have focused mostly on qualitative [30] and binary model responses. For instance, while [29] failed to detect racial or gender biases, their binary outcome measure may have been too coarse to facilitate detection. Ultimately, our approach allows us to measure disparities more accurately and with more variation, without the need to adopt subjective criteria.
Third, in an extension of previous efforts, we include 14 diverse domains that go beyond employment and are of particular salience and consequence (see section 3.1). Our approach also allows us to assess the sensitivity of biases to certain design features of the prompt. In effect, this allows us to explore mechanisms to counteract the existence of potential biases, leading to concrete and specific policy recommendations.
@@ 3 Methods and Design
We conduct a bias audit study of state-of-the-art LLMs. We emulate use cases across several domains in which language models could be used to give advice, taking into account different levels of context. Our approach involves receiving advice regarding a specific individual, and to vary that individual’s name. The names we choose are perceived to strongly correlate with race and gender, and we use direct model prompting as input to the models. We examine how these modifications affect the outputs of the models, focusing on eliciting quantitative responses for comparison. We adopt this design because probing the model directly with explicit mentions of race or gender can trigger mitigating measures taken by the developers. For instance, when specifying an individual’s race, GPT-4 will often refuse to respond or will provide responses that are otherwise insensitive to the remaining prompt. In addition, those deploying LLMs may take great care to blind the models to sensitive attributes, whereas our efforts are designed to surface implicit associations between race and less sensitive features that often evade censorship.
@@ 3.1 Prompt Design
For each scenario, we design several prompts following a structured process. These mutations are designed to identify bias, assess its heterogeneity, and explore potential mechanisms that may amplify or mitigate biases. We illustrate the design strategy with the example below. In addition, a summary of the different prompts is contained in Table 1.
@@
Scenario
Outcome
Variation
Context Level
Low
High
Numeric
@@
The first and perhaps most important aspect we vary is the name of the individual in each prompt. Varying the use of names between variations that are (perceived to be) strongly associated with a sensitive attribute like race or gender is a well-established practice in audit studies [3]. To enhance this methodology, we leverage findings from Gaddis’s research [9], which highlight that evidence is provided in less than 20% of studies that use names to signal race, frequently without exploring the effects of last names. Gaddis thus presents evidence of the racial perception associated with certain names and last names through a survey. We adopted the 40 names exhibiting the highest rates of congruent racial perception across race-gender groups. These names were paired with the last names with the highest percentage of Black and White individuals according to the U.S. Census Bureau (2012), consistent with those utilized in Gaddis’s study (namely, ”Washington” for Black individuals and ”Becker” for White individuals). The name ”Denzel Washington” was excluded to avoid association with the well-known actor. Overall, our list includes 14 names used in Bertrand and Mullainathan’s article [3], including one last name out of the two used in our design. In the example prompt, the name of the individual is “DaShawn Washington”, which is a name perceived to be strongly associated with the Black population in the U.S. A full list of names is contained in Table 3 in Appendix C.
@@
We measure the outcome quantitatively, rather than eliciting a qualitative description, as in [30, 29]. This is because a comparison of qualitative outputs requires a human, subjective assessment in order to produce comparability. In addition, the outcome we measure lies on a continuous scale or is measured in small discrete increments, such as the price in U.S. dollars or the probability of winning. In doing so, we depart from much of the existing literature, which often focuses on a binary assessment (e.g. “Should I make an offer to that job candidate? Yes/No”). We do this because a continuous measure allows for a more granular assessment of disparities, including measuring disparities more accurately and with more variation, without the need to adopt subjective criteria.
@@
In a last step, we vary more nuanced aspects within the scenario to illicit biases at a more granular level. For instance, in our “Sports” scenario, we assess both basketball as a sport with a high proportion of Black athletes, and Lacrosse, which has a historically low rate of Black athletes. In the above example, we consider the purchase of a bicycle. Other variations include the purchase of a car and of a house.
@@
Overall, we assess outcomes across 42 different prompt templates, all contained in Appendix A, across 40 names. The stochastic nature of language models can lead to variations in responses even under the same prompt. For that reason, we repeat our prompting for each combination of names and templates 100 times. The number of iterations was selected in an effort to balance both statistical power and costs. In total, our approach yields a comprehensive dataset of 168,000 responses. For 7 of these, the model output encompassed a range of values. In those instances, we chose the median value within the range. [^2] In 4 responses, the model did not provide a specific limit. For example, one response was ”…from around \$60,000 to over \$100,000 per year…” In that instance, we adopted \$109,000 as the upper limit, under the rationale that the model output would have included \$110,000 for greater limits. Overall, we were able to translate 99.96% of our responses directly into numeric values. For the remaining 0.04%, we imputed the median of the race/gender response in an effort to avoid missing values. This is because omitting missing values, while common, can induce significant bias [7]. Appendix D contains a more detailed explanation of our data post-processing methodology. We report the number of missing values in Table 4 in Appendix D.
@@ 3.2 Models
Our baseline model is OpenAI’s GPT-4, specifically the GPT-4-1106-preview variant. This version contains the latest knowledge up to April 2023. For consistency and to accurately reflect a potential use case of ChatGPT, we employ default parameters and system prompt across our evaluations. To assess our findings across LLMs, we incorporate additional proprietary models such as Google AI’s PaLM-2. To assess variation across model quality, we also compare the outcomes of GPT-4 to OpenAI’s GPT-3.5. For feasibility, we compare models only on the Purchase scenario, which is the scenario that exhibits the most significant differences across gender and race for GPT-4. These results are included in Appendix B.
@@ 4 Results
Figure 2 depicts the results of querying our baseline model (GPT-4) with prompts from the Purchase scenario. Without additional context, the model suggests a drastically higher initial offer when buying a bicycle/car from an individual whose name is perceived to be commonly held by white people. In contrast, names associated with the Black population in the U.S. receive substantially lower initial offers. Similarly, male-sounding names are associated with higher initial offers than female-sounding names.Unlike race-associations, the differences in offers for gender-associated names persist across all three variations.
As can be seen, these decrease rapidly when the model is provided with more detailed, qualitative information, although a statistically significant difference often remains. The exception to this general trend is the purchase of a house, where the provision of additional information induces racial biases and reverses gender biases. We hypothesize that this pattern might be the result of conditional disparities [^3] That is, conditioned on a particular type of home. exceeding unconditional disparities. For instance, in some of the responses to detailed queries about a home purchase, GPT-4 explicitly stated its assumption that the white person lives in a neighborhood with a higher price per square footage than the Black person. When providing the model with a numeric anchor, the responses become virtually identical across race and gender associations for all variations of the purchase prompt.
In Appendix B, we show that the results are substantively similar for Google’s PaLM-2 (Figure 5). Results are also qualitatively similar for GPT-3.5 (Figure 6). Importantly, the biases displayed by GPT-3.5 are not generally pronounced when compared with our results for GPT-4, suggesting that model quality is not a direct predictor of bias. Overall, the findings suggest that biases are prevalent across a variety of models, and are not limited to GPT-4.
@@
As can be seen, most scenarios display a form of bias that is disadvantageous to Black people and women. The only consistent exception to the rule is the basketball scenario. In it, consistent with our hypothesis, the model displays biases in favor of Black athletes. Overall, the results suggest that the model implicitly encodes common stereotypes, which in turn affects the model’s response. Because these stereotypes typically disadvantage the marginalized group, the advice given by the model does as well.
@@
As can be seen, the Black-perceived names yield systematically worse responses than white-perceived names. Similarly, female-perceived names yield systematically worse outcomes than male-perceived names. Overall, the findings suggest that the observed disparities are the result of a systematic bias, rather than a few outliers. In addition, Figure 4 suggests that intersectional biases are prevalent. In particular, Black, female-perceived names yield by far the worst response among all intersectional groups. Tables 5:9 in Appendix F provide disaggregated information on intersectionality for each scenario, variation and context level.
@@
Note: The figure shows the average standardized mean for each name, grouped by race and gender. This allows comparison despite different units of measurement in each scenario. Positions above or below the zero line suggest more or less favorable outcomes. We exclude all Sports scenarios since they were tailored to represent predominantly White or Black performance. See footnote 4.
Figure 4: Standardized Means for all Names.

### v3

# references.bib

@@ Abstract
We employ an audit design to investigate biases in state-of-the-art large language models, including GPT-4. In our study, we prompt the models for advice involving a named individual across a variety of scenarios, such as during car purchase negotiations or election outcome predictions. We find that the advice systematically disadvantages names that are commonly associated with racial minorities and women. Names associated with Black women receive the least advantageous outcomes. The biases are consistent across 42 prompt templates and several models, indicating a systemic issue rather than isolated incidents. While providing numerical, decision-relevant anchors in the prompt can successfully counteract the biases, qualitative details have inconsistent effects and may even increase disparities. Our findings underscore the importance of conducting audits at the point of LLM deployment and implementation to mitigate their potential for harm against marginalized communities.  [^1] Code is available at: https://github.com/AlexSalinas99/audit_llms.git
@@ 1 Introduction
Large Language Models (LLM) have dramatically surged in popularity over the recent years. Since the release of ChatGPT, LLMs - especially those with an accessible chat interface - have not only been used by experts, but are also becoming an increasingly common tool with significant benefits for laypeople. To that end, many commercial actors have already begun implementing LLMs in their operations, ranging from customer-facing chatbots to internal decision support systems [kanbach2023genai, constantz2023companies]. Additionally, users are turning into models to facilitate their day-to-day activities such as recruiting [ellis2024ai], negotiating [gold2024ai], or election forecasts [gujral2024llmshelppredictelections].
@@
The fairness of AI algorithms, including LLMs, has been a pernicious issue, motivating a growing literature and community of AI ethics research [CorbettDavies2023Fairness]. Disparities across gender and race, among other attributes, have especially preoccupied the field [caliskan2017semantics], leading to efforts to include bias auditing as an important component of AI harm mitigation in policy discussions and regulatory frameworks [Vecchione2021Algorithmic].
@@
Existing models have had relative success in mitigating biases arising from the explicit use of race or gender in the prompt. For instance, popular models like GPT-4 often refuse to provide an answer when prompted to produce information about a hypothetical individual when given that individual’s race. Similarly, companies that have access to sensitive features of their customers may simply foreclose access to this information from the LLM.
@@
However, biases can materialize not only through the explicit use of sensitive characteristics, but also by utilizing features that are (strongly) correlated with a person’s protected attributes. Mitigating the impact of such features can be more difficult, because, for one, their potential to cause disparate outcomes is often less salient, and for the other, these features may contain information that otherwise improves the utility of the model. In this study, we focus on an individual’s name as a feature of particular pertinence. Names strongly correlate with perceptions of race, raising the risk of creating significant disparities in model outputs, which can in turn harm marginalized communities. At the same time, in many practical applications, removing names might only come at a substantial cost. For instance, a chatbot that directly interacts with customers might significantly improve the experience via personalization if given access to the user’s name.
@@
We assess the name-sensitivity of the output produced by state-of-the-art language models. The names we choose are perceived to strongly correlate with race and/or gender, and we use direct model prompting as input to the models. Our assessment encompasses 42 idiosyncratic prompts. These prompts approximate use cases for 14 domains in which language models could be deployed to give advice to laypeople, such as in negotiations over the purchase of a car or in predicting election outcomes. We also assess how name-sensitivity interacts with the level of other useful information the model has access to in generating its output.
@@
We find significant disparities across names associated with race and gender in most scenarios we investigate, with varying effect sizes. The results are qualitatively similar across different models, including GPT-4o, GPT-4, GPT-3.5, Llama-3-70B, Mistral Large, and PaLM-2. Overall, we find that names associated with white men yield the most beneficial predictions, while those associated with Black women generate outcomes that disadvantage the individual in question. Providing the model with qualitative context about the person has an inconsistent effect on biases, at times amplifying and at times decreasing observed disparities; while a numeric anchor effectively removes name-based disparities in most scenarios we investigate. Our findings also suggest that the observed disparities are the result of a systematic bias, rather than the result of a few name outliers.
@@
Overall, the results suggest that the model implicitly encodes common stereotypes, which in turn affects the model response. Because these stereotypes typically disadvantage the marginalized group, the advice given by the model does as well. Our findings suggest name-based differences often materialize as disparities to the disadvantage of women, Black communities, and in particular Black women. The biases are consistent with common stereotypes prevalent in the U.S. population.
@@
The findings show that, despite efforts to mitigate biases and mount guardrails against disparate association with sensitive characteristics such as race and gender, LLMs still encode biases that translate into disparate outcomes. Despite earlier concerns over bias, even the latest models, such as GPT-4, are not immune to this problem. The findings raise concerns for companies that seek to incorporate LLMs into their operations, suggesting that masking race and gender may not be enough to prevent unwanted disparities. The findings also show that bias is pervasive and highlight the need for audits at the point of deployment and implementation, and not only at the development phase.
@@ 2.1 Defining Bias in Audit Studies
U.S. anti-discrimination laws generally encompass two distinct types of discriminatory conduct. First, there is “disparate treatment”, which refers to policies or actions that intentionally impose differential treatment due to protected characteristics like race or gender. The prototypical example of such policies are those that are explicitly conditioned on the protected characteristic. In effect, disparate treatment is often interpreted as corresponding to common, intuitive understandings of discrimination in which an individual receives a certain cost or benefit because of their race/gender. Disparate treatment by governmental actors is scrutinized and generally outlawed under the Fourteenth Amendment of the U.S. Constitution, and is similarly illegal in most decision-making by private actors due to the existence of several federal and state laws.
@@
[…]
@@ 2.2 Prior Literature
There is a substantial literature assessing bias in algorithms, including in medicine and health care [McCradden2020EthicalLimitations, Obermeyer2019RacialBias, Pfohl2021FairML, Goodman2018MachineLearning], law [Huq2019RacialEquity, Bent2019AlgorithmicAffirmative, Chander2016RacistAlgorithm, Kim2022RaceAwareAlgorithms, HoXiang2020AffirmativeAlgorithms, Gillis2022InputFallacy, YangDobbie2020EqualProtection, Mayson2019BiasInOut], and education [BakerHawn2022AlgorithmicBias, KizilcecLee2022AlgorithmicFairness]. The associated field is also referred to as “algorithmic fairness” [CorbettDavies2023Fairness], and its primary focus lies on assessing potential biases in algorithms that are used to assist human decision making. Researchers have also examined biases in automated speech recognition systems [Koenecke2020RacialDisparities] and facial recognition systems [khalil2020facialbias], among others.
[…]
@@
Audit studies have a long tradition in assessing biases in human decisions, going back to the civil rights movement [Vecchione2021Algorithmic]. Historically, they have involved pairs of ”testers” who go through the process of seeking benefits such as employment or housing. The pairs were made to look and behave similarly, with the main difference that a sensitive attribute–like race or gender–differs across the individuals in the pair. By measuring differences in outcomes, the researchers could identify biases in the decision making process of the entity under investigation (e.g., a housing corporation) as they relate to the sensitive attribute [yinger1998testing, Pager2007].
[…]
@@
[…]
Third, in an extension of previous efforts, we include 14 diverse domains that go beyond employment and are of particular salience and consequence (see section 3.1). Our approach also allows us to assess the sensitivity of biases to certain design features of the prompt.
@@ 3 Methods and Design
We conduct a bias audit study of state-of-the-art LLMs. We emulate use cases across several domains in which language models could be used to give advice, taking into account different levels of context. Our approach involves receiving advice regarding a specific individual, and varies that individual’s name. The names we choose are perceived to strongly correlate with race and gender, and we use direct model prompting as input to the models. We examine how these modifications affect the outputs of the models, focusing on eliciting quantitative responses for comparison. We adopt this design because probing the model directly with explicit mentions of race or gender can trigger mitigating measures taken by the developers. For instance, when specifying an individual’s race, GPT-4 will often refuse to respond or will provide responses that are otherwise insensitive to the remaining prompt. In addition, those deploying LLMs may take great care to blind the models to sensitive attributes, whereas our efforts are designed to surface implicit associations between race and less sensitive features that often evade censorship.
@@ 3.1 Prompt Design
For each scenario, we design several prompts following a structured process. These mutations are designed to identify bias, assess its heterogeneity, and explore potential mechanisms that may amplify or mitigate biases. We illustrate the design strategy with the example in Figure 1. In addition, a summary of the different prompts is contained in Table 1.
@@
Scenario Outcome Variation Context Level
Low High Numeric
@@
[…]
@@
[…]
@@
In a last step, we vary more nuanced aspects within the scenario to illicit biases at a more granular level. For instance, in our “Sports” scenario, we assess both basketball as a sport with a high proportion of Black athletes, and Lacrosse, which has a historically low rate of Black athletes. In Figure 1, we consider the purchase of a bicycle. Other variations include the purchase of a car and of a house.
@@
Overall, we assess outcomes across 42 different prompt templates (see Appendix A), across 40 names. The stochastic nature of language models can lead to variations in responses even under the same prompt. For that reason, we repeat our prompting for each combination of names and templates 100 times. The number of iterations was selected in an effort to balance both statistical power and costs. In total, our approach yields a comprehensive dataset of 168,000 responses. For 7 of these, the model output encompassed a range of values. In those instances, we chose the median value within the range. [^4] In 4 responses, the model did not provide a specific limit. For example, one response was ”…from around \$60,000 to over \$100,000 per year…” In that instance, we adopted \$109,000 as the upper limit, under the rationale that the model output would have included \$110,000 for greater limits. Overall, we were able to translate 99.96% of our responses directly into numeric values. For the remaining 0.04%, we imputed the median of the race/gender response in an effort to avoid missing values. This is because omitting missing values, while common, can induce significant bias [coppock2019avoiding]. Appendix D contains a more detailed explanation of our data post-processing methodology. We report the number of missing values in Table 4 in Appendix D.
@@ 3.2 Models
Our baseline model is OpenAI’s GPT-4, specifically the GPT-4-1106-preview variant. For consistency, and to accurately reflect a potential use case of ChatGPT, we employ default parameters and system prompts across our evaluations. To assess our findings across LLMs, we incorporate additional proprietary models such as Google AI’s PaLM-2 and Mistral-Large, as well as open-source models such as Llama-3 70B. To assess variation across model quality, we also compare the outcomes of GPT-4 to OpenAI’s GPT-3.5 and GPT-4o.
@@ 4 Results
Figure 2 depicts the results of querying our baseline model (GPT-4) with prompts from the Purchase scenario. Without additional context, the model suggests a drastically higher initial offer when buying a bicycle/car from an individual whose name is perceived to be commonly held by white people. In contrast, names associated with the Black population in the U.S. receive substantially lower initial offers. Similarly, male-associated names are associated with higher initial offers than female-associated names. Unlike race-associations, the differences in offers for gender-associated names persist across all three variations.
As can be seen, these biases decrease when the model is provided with more detailed, qualitative information, although a statistically significant difference often remains. The exception to this general trend is the purchase of a house, where the provision of additional information induces racial biases and reverses gender biases. We hypothesize that this pattern might be the result of conditional disparities [^5] That is, conditioned on a particular type of home. exceeding unconditional disparities. For instance, in some of the responses to detailed queries about a home purchase, GPT-4 explicitly stated its assumption that the white person lives in a neighborhood with a higher price per square footage than the Black person. When providing the model with a numeric anchor, the responses become virtually identical across race and gender associations for all variations of the purchase prompt.
In Appendix B, we show that the results are substantively similar for all tested models, which include Google’s PaLM-2 (Figure 5), GPT-3.5 (Figure 6), GPT-4o (Figure 7), Llama-3 70B (Figure 8), and Mistral Large (Figure 9). Importantly, the biases displayed by GPT-3.5 are not generally pronounced when compared with our results for GPT-4, suggesting that model quality is not a direct predictor of bias. Overall, the findings suggest that biases are prevalent across a variety of models, and are not limited to GPT-4.
@@
As can be seen, most scenarios display a form of bias that is disadvantageous to Black people and women. The only consistent exception to this pattern is the basketball scenario. In it, consistent with our hypothesis, the model displays biases in favor of Black athletes. Overall, the results suggest that the model implicitly encodes common stereotypes, which in turn affects the model’s response. Because these stereotypes typically disadvantage the marginalized group, the advice given by the model does as well.
@@
As can be seen, the Black-perceived names yield systematically worse responses than white-perceived names. Similarly, female-perceived names yield systematically worse outcomes than male-perceived names. Overall, the findings suggest that the observed disparities are the result of a systematic bias, rather than a few outliers. Next, we examine biases for Black/white-associated, male/female-associated names separately. In doing so, we note that this analysis does not equate to an evaluation of intersectionality, as different identities may co-construct in ways that are not captured by the interaction of race and gender [intersect, factoring]. At the same time, we believe that this analysis can offer valuable insights into bias directed against individuals who the model perceives to be associated with multiple minority identities, and thus may be particularly vulnerable. Figure 4 suggests Black, female-perceived names yield by far the worst response among all minority groups. Tables 5:9 in Appendix F provide disaggregated information on this result for each scenario, variation and context level. Furthermore, our findings in Figures 15 and 16 in Appendix H reveal that these biases are pervasive across all models we audited. Notably, the other models examined exhibited even greater biases than GPT-4, suggesting that these issues could be more severe across the broader landscape of large language model usage.
@@
Note: The figure shows the average standardized mean for each name, grouped by race and gender. This allows comparison despite different units of measurement in each scenario. Positions above or below the zero line suggest more or less favorable outcomes. We exclude all Sports scenarios since they were tailored to represent predominantly White or Black performance. See footnote 6.